# A modern perspective on rational homotopy theory

### v1

@@ 11.1. The model category of topological spaces
    It is not combinatorial because the category ${\mathcal{T}\mathrm{op}}$ is not locally presentable [AR94, Example 1.18(5)].
The latter difference raises the question of whether ${\mathcal{T}\mathrm{op}}$ can be left Bousfield localized. To that end, we remark that ${\mathcal{T}\mathrm{op}}$ shares with ${\mathcal{SS}\mathrm{ets}}$ the feature of being a cellular model category [Hir03, Proposition 12.1.4]. This property is a different way of strengthening the property of being a cofibrantly generated model category, and it allows us to left Bousfield localize ${\mathcal{T}\mathrm{op}}$, for Theorem 7.3 holds true if we replace “combinatorial” with “cellular” throughout [Hir03, Theorem 4.1.1].
11.2. The family of model structures on topological spaces
Our work in Section 9 yields the promised family of model structures on topological spaces.
Theorem 11.7.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a model structure $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{T}\mathrm{op}}$ on the category ${\mathcal{T}\mathrm{op}}$ of topological spaces in which a map $f\colon X\rightarrow Y$ is
  1. (1)
    a weak equivalence if and only if, for every $M$-local space $K$, the map
    $f^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,K)$
    is a weak homotopy equivalence; and
@@ 11.2. The family of model structures on topological spaces
Proof.
Proposition 9.7 informs us that $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{T}\mathrm{op}}$ is the left Bousfield localization of ${\mathcal{T}\mathrm{op}}$ at the set of maps
$\left\{S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}\mid n\geq 2\right\}.$
@@
$|B_{M^{-1}\mathbb{Z}}|=\left\{|\operatorname{Sing}(S^{n})|\rightarrow|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|\mid n\geq 2\right\}.$
Thus, by [Hir03, Theorem 3.3.20], the Quillen equivalence
$|-|\colon\mathcal{SS}\mathrm{ets}\rightleftarrows\mathcal{T}\mathrm{op}\colon\operatorname{Sing}$
induces a Quillen equivalence
@@
This Quillen equivalence captures the fact that the model categories $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}$ and $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{T}\mathrm{op}$ both describe the $M$-local homotopy theory beyond $1$-connected spaces from Section 9.
Lastly, we distinguish the model category corresponding to the case $M^{-1}\mathbb{Z}=\mathbb{Q}$.
Theorem 11.9.
There is a model structure $\mathcal{L}_{\mathbb{Q}}{\mathcal{T}\mathrm{op}}$ on the category ${\mathcal{T}\mathrm{op}}$ of topological spaces in which a map $f\colon X\rightarrow Y$ is
  1. (1)
    a weak equivalence if and only if, for every rational space $K$, the map
    $f^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,K)$
    is a weak homotopy equivalence; and
  2. (2)
A space $X$ is fibrant in $\mathcal{L}_{\mathbb{Q}}{\mathcal{T}\mathrm{op}}$ if and only if $X$ is rational. The model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{T}\mathrm{op}}$ is cellular, left proper, and simplicial with the structure from Proposition 11.5.
12. Rational homology versus rational homotopy
We conclude this paper by contrasting Bousfield’s localization of spaces with respect to rational homology to our localization of spaces with respect to rational homotopy. Even though rational homology and rational homotopy agree for $1$-connected spaces, this agreement does not extend to all spaces. Therefore, we verify that the two localized model structures on spaces are not equal.
12.1. Localization of topological spaces with respect to homology
Our argument is best presented using topological spaces, so we formulate the analog of Theorem 8.7 for the category $\mathcal{T}\mathrm{op}$ of topological spaces.
Theorem 12.1 ([Bou75, Theorem 10.2]).
@@ 12.1. Localization of topological spaces with respect to homology
By [Bou97, §2.5], $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{T}\mathrm{op}$ is the left Bousfield localization of $\mathcal{T}\mathrm{op}$ at a single map, so it inherits the properties of being cellular, left proper, and simplicial.
When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we get a model structure whose weak equivalences are the rational homology isomorphisms.
Theorem 12.3 ([Bou75, Theorem 10.2]).
There is a model structure $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$ on the category of topological spaces in which a map $f$ is
  1. (1)
    a weak equivalence if and only if $f$ is a rational homology isomorphism; and
  2. (2)
12.2. Rational homology versus rational homotopy
We show that Bousfield’s localization $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$ with respect to rational homology does not coincide with our localization $\mathcal{L}_{\mathbb{Q}}\mathcal{T}\mathrm{op}$ with respect to rational homotopy from Theorem 11.9. This result stems from the fact that rational homology and rational homotopy cease to agree beyond the realm of $1$-connected spaces, and our proof comprises a particular instance of this phenomenon.
@@ 12.2. Rational homology versus rational homotopy
We show that the unique map
$f\colon\mathbb{R}P^{2}\rightarrow|\Delta[0]|$
is a weak equivalence in $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$ but not a weak equivalence in $\mathcal{L}_{\mathbb{Q}}\mathcal{T}\mathrm{op}$.
Firstly, $f$ is a weak equivalence in $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$ because, by Example 8.3, $f$ is a rational homology isomorphism.
To show that $f$ is not a weak equivalence in $\mathcal{L}_{\mathbb{Q}}\mathcal{T}\mathrm{op}$, we recall from Example 9.3 that $\mathbb{R}P^{\infty}$ is a rational space because its higher homotopy groups vanish. This recollection renders it sufficient to show that
$f^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}(|\Delta[0]|,\mathbb{R}P^{\infty})\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(\mathbb{R}P^{2},\mathbb{R}P^{\infty})$
is not a weak homotopy equivalence. In particular, it suffices to show that the induced map on path components
$f^{\ast}\colon[|\Delta[0]|,\mathbb{R}P^{\infty}]\rightarrow[\mathbb{R}P^{2},\mathbb{R}P^{\infty}]$
is not a bijection. To that end, we find a map from $\mathbb{R}P^{2}$ to $\mathbb{R}P^{\infty}$ that is not null. Such a map is the inclusion
$i\colon\mathbb{R}P^{2}\rightarrow\mathbb{R}P^{\infty},$
for its induced map on cohomology rings with $\mathbb{Z}/2\mathbb{Z}$-coefficients is the quotient map
$i^{\ast}\colon(\mathbb{Z}/2\mathbb{Z})[x]\rightarrow(\mathbb{Z}/2\mathbb{Z})[x]/(x^{3}),$
which is not trivial. Overall, $f$ is a weak equivalence in $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$ but not a weak equivalence in $\mathcal{L}_{\mathbb{Q}}\mathcal{T}\mathrm{op}$. ∎
Thanks to Example 3.2 and Theorem 4.1(2), we can safely apply the singular complex functor throughout our proof of Proposition 12.4 to obtain the analogous result for simplicial sets.
Proposition 12.5.
The model structures $\mathcal{L}_{H\mathbb{Q}}\mathcal{SS}\mathrm{ets}$ and $\mathcal{L}_{\mathbb{Q}}\mathcal{SS}\mathrm{ets}$ on simplicial sets are not equal.

### v2

@@ 10.3. The family of model structures on topological spaces
In particular, we obtain a model category whose weak equivalences are the rational homotopy equivalences and whose fibrant objects are the rational spaces.
@@
$|T_{M^{-1}\mathbb{Z}}|=\{|\operatorname{Sing}(S^{n})|\rightarrow|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|\mid n\geq 2\}.$
Thus, by [26, Theorem 3.3.20], the Quillen equivalence $|-|\colon\mathcal{SS}\mathrm{ets}\rightleftarrows\mathcal{T}\mathrm{op}\colon\operatorname{Sing}$ induces a Quillen equivalence
@@
10.4. Rational homotopy versus rational homology
We conclude this paper by contrasting our localization of all spaces at rational homotopy to Bousfield’s localization at rational homology. Our argument is best presented using topological spaces, so we first formulate Theorem 8.7 for topological spaces.
Theorem 10.12 ([7, Theorem 10.2]).
@@ 10.4. Rational homotopy versus rational homology
By [9, §2.5], $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{T}\mathrm{op}$ is the left Bousfield localization of $\mathcal{T}\mathrm{op}$ at a single map, so it inherits the properties of being cellular, left proper, and simplicial.
We show that our localization at rational homotopy $\mathcal{L}_{\mathbb{Q}}\mathcal{T}\mathrm{op}$ differs from that at rational homology $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$.
@@
We show that $c\colon\mathbb{R}P^{2}\rightarrow|\Delta[0]|$ is a weak equivalence in $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$ but not in $\mathcal{L}_{\mathbb{Q}}\mathcal{T}\mathrm{op}$. Firstly, $c$ is a weak equivalence in $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$ because, by Example 8.3, $c$ is a rational homology isomorphism. However, $c$ is not a weak equivalence in $\mathcal{L}_{\mathbb{Q}}\mathcal{T}\mathrm{op}$ because, by Example 9.3, $c$ is not a rational homotopy equivalence. ∎